OpenStudy (anonymous):

Rewrite with positive exponents. Assume that even roots are of nonnegative quantities and that all denominators are nonzero. (6xy/y)^-3

OpenStudy (anonymous):

3x + 24y = 24, you could start by isolating x in this equation

OpenStudy (anonymous):

so, 9(3x+24y)=24*9 and then just add 27x+216y=216 +27x-15y=-15 -------------- -15y=216-15 -15y=201 --------------- 15 15 y=13.4 is this right

OpenStudy (anonymous):

no were simply rearranging this equation to make in x= format 3x + 24y = 24 subtract the 24y on both sides 3x=24-24y then divide both sides by 3 x=24/3 - 24y/3 reduce it x=8-8y

OpenStudy (anonymous):

do u understand this part?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so now all we do is substitute that into the second equation 27(8-8y) – 15y = -15

OpenStudy (anonymous):

216-216y-15y=-15 216-231y=-15 subtract 216 from both sides -231y=-15-216 -231y=-231 divide y=-y is this right

OpenStudy (anonymous):

your correct up until here -231y=-231 divide both sides by -231 y=1

OpenStudy (anonymous):

oh ic thanks

OpenStudy (anonymous):

is that the end of the problem

OpenStudy (anonymous):

no we just need to find the x value now

OpenStudy (anonymous):

all you have to do is replace y with 1 on either of those equations then solve for x

OpenStudy (anonymous):

3x+24*1=24 3x+24=24 ok idk

OpenStudy (anonymous):

u got it right so far keep going

OpenStudy (anonymous):

3x+24=24 subtract 24 from both sides 3x=24-24

OpenStudy (anonymous):

or would i add 24

OpenStudy (anonymous):

ya u would subtract the 24

OpenStudy (anonymous):

but 3x=24-24 =3x=0 and then divide both sides and get x=0 is that right

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so is that it then is the problem done

OpenStudy (anonymous):

with the answer being (x,y)=(1,0)

OpenStudy (anonymous):

thank you

OpenStudy (unklerhaukus):

check you answer 3x + 24y = 24 27x – 15y = -15 x=1 y=0 3 ≠ 24 27≠ -15

OpenStudy (anonymous):

oh ya you had the answer right you just wrote it wrong

OpenStudy (unklerhaukus):

3x + 24y = 24 27x – 15y = -15 x=0 y=1 24 = 24 – 15= -15